A cheetah accelerates from rest at 9.2 m/s². How long will it take the cheetah to reach a distracted gazelle that is 66 m away from the cheetah?


1
Expert's answer
2021-09-22T07:06:05-0400

Using the kinematic formula


"d = \dfrac{at^2}{2}"

where "d=66m" is the distance, "a = 9.2m\/s^2" is the acceleration, and "t" is the time, find:


"t = \sqrt{\dfrac{2d}{a}}\\
t = \sqrt{\dfrac{2\cdot 66m}{9.2m/s^2}} \approx 3.8 s"

Answer. 3.8 s.